OpenStudy (anonymous):

Circles O and P are tangent and have radii of 16 and 4, respectively. Find the length of their common tangent.

OpenStudy (agent0smith):



OpenStudy (mathstudent55):

 \[(r - s)^{2} + x ^{2} = (r + s)^{2}\] \[x ^{2} = (r + s)^{2} - (r - s)^{2}\] \[x ^{2} = r ^{2} + 2rs + s ^{2} - r ^{2} + 2rs - s ^{2}\] \[x ^{2} = 4rs\] \[x = 2\sqrt{rs}\]